OpenStudy (needhelpp101):

Provide an appropriate response. Write the next equation in the pattern. 2 - 1 = 1. 3 - 2 = 1. 5 - 3 = 2. 8 - 5 = 3 A. 13 - 8 = 5 B. 10 - 5 = 5 C. 11 - 3 = 8 D. 13 - 5 = 8

OpenStudy (anonymous):

its b. All pattern based vertically.

OpenStudy (anonymous):

2+3=5, 3+5=8, 5+8=13 1+2=3, 2+3=5, 3+5=8 and so on..

OpenStudy (anonymous):

sorry I meant that its A

OpenStudy (needhelpp101):

Thanks Heycoa :)